OpenStudy (anonymous):

What is the rate of decay of awesome-10 which has a half-life of 650 days?

OpenStudy (anonymous):

I can't figure out how to solve these rate of decay questions, so I just made up my own as a way to help me learn this. How would I solve this?

OpenStudy (anonymous):

if it has a half life of 650 time units it means every 650 time units 1/2 of the samle dissapears.

OpenStudy (anonymous):

Yep. :) How would I find the rate of decay, though?

OpenStudy (anonymous):

halflife= -ln2/k that is the half life equation

OpenStudy (anonymous):

I have a formula \[N=N _{0}e ^{kt}\] How/where would I plug everything in?

OpenStudy (anonymous):

I'm not sure I understand @Spectrum

OpenStudy (anonymous):

I know spectrum

OpenStudy (anonymous):

i gave him all right answers

OpenStudy (anonymous):

for his teset he did good to

OpenStudy (anonymous):

I think they are asking for k you got N=1/2No t=650 so you plug all numbers in and find k

OpenStudy (anonymous):

I mean, I don't understand what he meant. I'm having a lot of trouble with logarithms, so I'm not quick on my feet. :P

OpenStudy (anonymous):

\[N=\frac{ 1 }{ 2 }N _{0}^{650k}\] Okay, so what do I do next?

OpenStudy (anonymous):

Would I put some test number in for No? Like 100 or something?

OpenStudy (anonymous):

no, you plug 1/2No for Nthen you can simplify for No